# Photon production in top quark events at ATLAS and CMS

### v1

@@ 1 Introduction
With the data collected at the LHC in its Run 2, the $\mathrm{t\bar{t}\gamma}$ process entered the precision era, with inclusive cross sections measured with a precision down to $\sim$4%. Differential measurements have also been performed in several final states, by both ATLAS and CMS collaborations. For the $\mathrm{tq\gamma}$ process, observation (evidence) was reported by ATLAS (CMS), and inclusive cross section measurements have been performed. The $\mathrm{tW\gamma}$ process, which is an important background to $\mathrm{t\bar{t}\gamma}$ and interferes with it beyond the leading order, has only been measured summed with $\mathrm{t\bar{t}\gamma}$ by ATLAS in $\mathrm{e\mu}$ final states.
@@ 2 Theoretical and experimental challenges
The $\mathrm{t\bar{t}\gamma}$ process includes events with photons emitted from the initial state quarks, the top quarks, and the top quark decay products. These are experimentally indistinguishable, though their kinematics differ enough that it is possible to create phase spaces enriched in one photon origin with respect to the others. For example, high $p_{T}$ photons originate mostly from the initial state or the top quarks. Nevertheless, in a $\mathrm{t\bar{t}\gamma}$ analysis, one needs to model all possible photon origins. At leading order (LO) in quantum chromodynamics (QCD), this can be done relatively easily, however, at next-to-LO (NLO), it becomes computationally unfeasible. One possible approach is to use a LO model and scale its cross section using a so-called k-factor, and another is to use an NLO model which includes only photons from the initial state or off-shell top quarks (referred to as "$\mathrm{t\bar{t}\gamma}$ production") stitched with a LO model filtered to include only photons from the on-shell tops or their decays (called "$\mathrm{t\bar{t}\gamma}$ decay" in the following).
An additional challenge comes from the modelling of $\mathrm{tW\gamma}$, which interferes with $\mathrm{t\bar{t}\gamma}$ beyond the LO. This interference can in principle be removed using a diagram removal strategy [5], however published results so far only use $\mathrm{tW\gamma}$ simulated samples at LO, where this interference is not present.
@@
In CMS, a strategy known as the ABCD method is typically used, where control regions enhanced in fake photons are built by inverting the selections on the charged isolation of the photon candidates and/or the width of the corresponding electromagnetic shower. The rate of fake photons is then determined in these regions and transported to the signal region. In the ATLAS analyses, a similar strategy is employed to estimate the contribution of photons from hadrons and pileup, while misreconstructed electrons are estimated from the fraction of electron–positron candidates from $\mathrm{Z\to ee}$ decays that are reconstructed as $\mathrm{Z\to e\gamma}$.
@@ 3 Inclusive cross section measurements of $\mathrm{t\bar{t}\gamma}$
$\sigma_{\mathrm{t\bar{t}\gamma}\text{ (production)}}=322\pm 5\text{ (stat)}\pm 15\text{ (syst)}~{}\mathrm{fb}\,\,(5\%)\,,$
in agreement with the prediction of $299\pm 31$ fb from MadGraph5_aMC@NLO. The measurement is dominated by the systematic uncertainties, mainly those on the modelling of $\mathrm{t\bar{t}\gamma}$, the normalization of the $\mathrm{t\bar{t}\gamma}$ decay component, as well as jet and b tagging uncertainties.
The CMS measurements focus on the lepton+jets [7] and dilepton [8] channels separately, and measure the fiducial cross sections for the total $\mathrm{t\bar{t}\gamma}$ process (production+decay). In this case, the whole signal is modelled at LO in QCD with MadGraph5, scaled to NLO using k-factors. Once again the nonprompt photon contribution is estimated with data-driven methods. Figure 1 (right) shows a comparison between data and simulation for photon $p_{T}$ in the dilepton channel. In the lepton+jets channel, a simultaneous fit to two signal regions and six control regions is performed, while in the dilepton channel, one single signal region is fitted. The main systematic uncertainties affecting the measurement are those on signal modelling, background normalization, the estimation of nonprompt photons, and the luminosity. All results are compatible with theoretical calculations performed at NLO in QCD.
@@ 6 Top quark charge asymmetry in $\mathrm{t\bar{t}\gamma}$ events
The top quark charge asymmetry ($A_{c}$) is an anisotropy in the angular distributions of the final-state top quark and antiquark, $A_{C}=\frac{\sigma_{+}-\sigma_{-}}{\sigma_{+}+\sigma_{-}}$, where $\sigma_{+(-)}$ is the cross section for positive (negative) values of $|y(\mathrm{t})|-|y(\mathrm{\bar{t}})|$. For $\mathrm{t\bar{t}}$ production, the SM prediction at NLO in QCD is of about 0.6%. In $\mathrm{t\bar{t}\gamma}$ events, the charge asymmetry is potentially enhanced and expected to have the opposite sign compared to $\mathrm{t\bar{t}}$, due to photon emission diagrams contributing to the interference, already at LO. The SM prediction at NLO varies between [-0.5%,-2%], depending on phase space choices.

### v2

@@ 1 Introduction
With the data collected at the LHC during Run 2, the $\mathrm{t\bar{t}\gamma}$ process entered the precision era, with inclusive cross sections measured with a precision down to $\sim$4[8]. Differential measurements have also been performed in several final states, by both ATLAS and CMS collaborations [6, 7, 8]. For the $\mathrm{tq\gamma}$ process, observation (evidence) was reported by ATLAS (CMS) [11, 12], and inclusive cross section measurements have been performed. The $\mathrm{tW\gamma}$ process, which is an important background to $\mathrm{t\bar{t}\gamma}$ and interferes with it beyond the leading order, has only been measured summed with $\mathrm{t\bar{t}\gamma}$ by ATLAS in $\mathrm{e\mu}$ final states [9].
@@ 2 Theoretical and experimental challenges
The $\mathrm{t\bar{t}\gamma}$ process includes events with photons emitted from the initial state quarks, the top quarks, and the top quark decay products. These are experimentally indistinguishable, though their kinematics differ enough that it is possible to create phase spaces enriched in photons from a specific origin. For example, high $p_{T}$ photons originate mostly from the initial state or the top quarks. Nevertheless, in a $\mathrm{t\bar{t}\gamma}$ analysis, one needs to model all possible photon origins. At leading order (LO) in quantum chromodynamics (QCD) this can be done relatively easily, however, at next-to-LO (NLO), simulating the full $2\to 7$ process is not possible. One possible approach is to use a LO model and correct its cross section to the NLO value, and another is to use an NLO model which includes only photons from the initial state or off-shell top quarks (referred to as "$\mathrm{t\bar{t}\gamma}$ production") stitched with a LO model filtered to include only photons from the on-shell tops or their decays (called "$\mathrm{t\bar{t}\gamma}$ decay" in the following).
An additional challenge comes from the modelling of $\mathrm{tW\gamma}$, which interferes with $\mathrm{t\bar{t}\gamma}$ beyond the LO. This interference can in principle be removed using diagram removal (DR) or diagram subtraction (DS) strategies [5], however published results so far only use $\mathrm{tW\gamma}$ simulated samples at LO, where this interference is not present.
@@
In CMS, a strategy known as the ABCD method is typically used [13], where control regions enhanced in fake photons are built by inverting the selections on the charged isolation of the photon candidates and/or the width of the corresponding electromagnetic shower. The rate of fake photons is then determined in these regions and transported to the signal region. In the ATLAS analyses, a similar strategy is employed to estimate the contribution of photons from hadrons and pileup, while misreconstructed electrons are estimated from the fraction of electron–positron candidates from $\mathrm{Z\to ee}$ decays that are reconstructed as $\mathrm{Z\to e\gamma}$.
@@ 3 Inclusive cross section measurements of $\mathrm{t\bar{t}\gamma}$
$\sigma_{\mathrm{t\bar{t}\gamma}\text{ (production)}}=319\pm 4\text{ (stat)}\pm 15\text{ (syst)}~{}\mathrm{fb}\,\,(5\%)\,,$
in agreement with the prediction of $296\pm 30$ fb from MadGraph5_aMC@NLO. The measurement is dominated by the systematic uncertainties, mainly those on the modelling of $\mathrm{t\bar{t}\gamma}$, the normalization of the $\mathrm{t\bar{t}\gamma}$ decay component, as well as jet and b tagging uncertainties.
The CMS measurements focus on the lepton+jets [7] and dilepton [8] channels separately, and measure the fiducial cross sections for the total $\mathrm{t\bar{t}\gamma}$ process (production+decay). In this case, the whole signal is modelled at LO in QCD with MadGraph5, scaled to NLO. Once again the nonprompt photon contribution is estimated with data-driven methods. Figure 1 (right) shows a comparison between data and simulation for photon $p_{T}$ in the dilepton channel. In the lepton+jets channel, a simultaneous fit to several signal and control regions is performed, while in the dilepton channel, one single signal region is fitted. The main systematic uncertainties affecting the measurement are those on signal modelling, background normalization, the estimation of nonprompt photons, and the luminosity. All results are compatible with theoretical calculations performed at NLO in QCD.
@@ 6 Top quark charge asymmetry in $\mathrm{t\bar{t}\gamma}$ events
The top quark charge asymmetry ($A_{c}$) is an anisotropy in the angular distributions of the final-state top quark and antiquark, $A_{C}=\frac{\sigma_{+}-\sigma_{-}}{\sigma_{+}+\sigma_{-}}$, where $\sigma_{+(-)}$ is the cross section for positive (negative) values of $|y(\mathrm{t})|-|y(\mathrm{\bar{t}})|$. For $\mathrm{t\bar{t}}$ production, the SM prediction at NLO in QCD is of about 0.6%. In $\mathrm{t\bar{t}\gamma}$ events, the charge asymmetry is potentially enhanced and expected to have the opposite sign compared to $\mathrm{t\bar{t}}$, due to photon emission diagrams contributing to the interference, already at LO. The SM prediction at NLO varies between [-0.5%,-2%], depending on phase space choice.